The work function of a photoelectric material is 4.0 eV. (a) What is the threshold wavelength? (b) Find the wavelength of light for which the stopping potential is 2.5 V.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. ∅ = 4 eV = 4 * 1.6 *10^-19 J a) Threshold wavelength = λ ∅ = hcλ ⇒ λ = hC/∅ = 6.63 * 10^-34 *3 * 10^8/4 * 1.6 *10^-19 = 6.63 * 3/6.4 * 10^-27/10^-9 = 3.1 * 10^7 m = 310 nm. b) Stopping potential is 2.5 V E = ∅ + eV ⇒ hcλ = 4 * 1.6 * 10^-19 + 1.6 * 10^-19 * 2.5 ⇒ λ = 6.63 *10^-34 *3 * 10^8/λ * 1.6 * 10^-19 = 4 + 2.5 ⇒ 6.63 * 3 * 10^-26/1.6 * 10^-19 * 2.5 = 1.9125 * 10^-7 = 190 nm.